OpenStudy (anonymous):

Write each expression as a single logarithm. Assume that all the variables represent positive numbers.

OpenStudy (anonymous):

\[\ln _{2} x ^{2} - \ln _{2}xy +\ln _{2}y^{2}\]

OpenStudy (sweetburger):

are those base 2 or base 3 i cant see

OpenStudy (anonymous):

the bases are 2s

OpenStudy (anonymous):

hi how are you?

OpenStudy (luigi0210):

Rule: \[\LARGE \log_{x}a-\log_{x}b=\log_{x}\frac{a}{b}\]

OpenStudy (anonymous):

Yes

OpenStudy (luigi0210):

Do you know the rules?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif use rule 2 listed there, and then rule 1

OpenStudy (luigi0210):

So what seems to be the problem.

OpenStudy (anonymous):

 correct?

OpenStudy (anonymous):

Can i ask another question? Sorry - that was a stupid one.

OpenStudy (luigi0210):

Sure~

OpenStudy (anonymous):

\[\log _{2}x = 2\log _{2}7 + \log _{2}5\]